20140320

MATGUI command to start the Mat camera on top of windbracing and check for laser pointing on wide FoV. Username: ???, pwd: ???

20140322

6:00 - II test of new collimation tables. JH wants to check the preset acquire accuracy with and without the dichroic in.

Telescope pointed to BS9143, mr=14, 62 elev.

Preset acquire (agw on axis) in parallactic mode, position angle 90deg (PE90). The rotator angle is at 84.8deg: not good, JH prefers having PE0.

6:20 - New preset acquire with PE0 on the same target, rotator angle is at 3.39deg.

Preset active and guiding on.

Dichroic slided in and new preset acquire authorizing LUCI@ARGOS focal station: we find the star within 3": good.

6:35 - dichroic slided out and preset acquire authorizing LUCI@RFBG: we find the star within 1": the test of the new collimation table made by JH is succesful!

6:50 - Handover to Laser operation, spotters outside and shutter test performed.

6:57 - Start of laser propagation, laser pointing adjustment with MAT, cirrus at 10km make laser beams blobby. On PatCams we see M2 spider. TN: 20140322_070531

13444222674_46701c6009_o.jpg patrol_camera_blue.jpeg

Loaded reconstructor for PatCam vs LM1 offload: 20140320_063219, 3 manual iterations.

TN: 20140322_071605 spot are closer as possible to LGSW FoV.

We try to acquire the IM between the PatCam and the LAS periscopes, other adjustment with LM1.

20140322_072523: PatCam to periscope IM with 5arcsec movement, reconstructor loaded.

We try to offset the LGS position error to periscopes.

TN: 20140322_072847 fit is not valid!!!

RED PatCam offset evaluated by hand and periscopes moved by hand: the spot is properly centered on PatCam but we see no light on PnCCD (just the Blue pupil).

TN: 20140322_074331 3 LGS recentered

LAN focus stage set to -100k steps

LM1 moved to recenter just BLUE pupil on PnCCD

Time of flight test: 90us better than 70us

Test of ToF sweep between 9 and 16km: too fast the change of ToF. RED and YELLOW pupils change behaviour with sweep. BLUE one stay fixed. We try also a hand sweep. 90km seems the better position.

We sweep also the LAN focus. We put it to -60k.

TN: 20140322_080643

We close both PI and pupil loops

TN: 20140322_080832

20140322_080832_lgsw_ccdframes_averaged.jpg

TN: 20140322_081238 just pupil loop closed

20140322_081238_lgsw_ccdframes_averaged.jpg

We try to record the IM between PI and periscope: 20140322_082121

The measure offset are too big. We try a IMX0.1

8:38 - periscope gain at 0.9, bad. gain at 0.6. The fit procedure is not working when the spot is in the LGSW FoV. A quad cell algorithm could be implemented.

8:43 - Invalid fit fixed by LB putting 0 to the measured position in case of non converging fit...

8:46 - We try to close the pupil and jitter loops

8:56 - Sweeo ToF: 90us choosed by eye, this corresponds to 13.5 km

8:58 - We try to apply a static shape to M2: 2um iof Z5

TN: 20140322_090131

New slope_null with M2 flat: 20140322_090256

Then we apply 10um of Z5 to M2

9:02 - we try to seep the LAN focus taking snapshots to measure the FHWM:

TN: 20140322_091455 LAN @ -180k steps, not good

TN: 20140322_091527 LAN @ -160k

TN: 20140322_091552 LAN @ -140k

lgsw_ccdframes_averaged_160k-140k.jpg

TN: 20140322_091827 LAN @ -120k

TN: 20140322_091859 LAN @ -100k

TN: 20140322_091936 LAN @ -80k

TN: 20140322_092105 LAN @ -60k, we recenter the LGS, not good

TN: 20140322_092133 LAN @ -60k

TN: 20140322_092209 LAN @ -40k

We recenter the LGS

TN: 20140322_092408 LAN @ -20k

Now we put LGS on PatCams and we sweep again:

TN: 20140322_094639 LAN @ -180k steps, spot on PatCams

TN: 20140322_094734 LAN @ -160k steps, spot on PatCams

TN: 20140322_094757 LAN @ -140k steps, spot on PatCams

TN: 20140322_094839 LAN @ -120k steps, spot on PatCams

TN: 20140322_094931 LAN @ -100k steps, spot on PatCams

TN: 20140322_095009 LAN @ -80k steps, spot on PatCams

TN: 20140322_095055 LAN @ -60k steps, spot on PatCams

Yellow LGS appears a little more out of focus with respect the other 2 LGS...

The roundish spot is at -120k.

Recenter LGS with LM1 and periscope loops.

TN: 20140322_102601 mix of internal source + LGS, we're vignetting the red LGS

11:39 - Switched on FLAO, telescope is collimated on WT10_292 (mr=6.5, 14.5deg from Zenith), AO loop closed and new flat saved for M2 (ARGOS and ARGOS1)

New slope_null with flattened M2: 20140322_120511

12:31 - Telescope back to Zenith having set ARGOS1 flat, M2 is back to banana shape...

TN: 20140322_125803 good snapshot to define subapertures.

20140323

Defined a new target position for PatCams: 20140322_205100

CalUnitSystem light on PatCams in DayTime, TN: 20130322_220800

1:20 - Test of PockelsCells: 100ns opening time counts: 1900, 900, 280

200ns: 2050, 950, 300

TN: 20140323_012459 solar ligth on M1 -> good illumination to measure FoV diameter

We put diodes to 0.2A and we scan the Pockels cells opening time: 100ns: 9000, 4000, 1600

500ns: 10200, 4800, 2000

1000ns: 11000, 6000, 2600

1500ns: 12800, 7200, 3300

3:16 - Preset acquire authorizing LUCI@RFBG on ACT0284 (about 4deg from M35, EL 60deg), dichroic slided in, the target is offsetted by [6.0; 1.0]arcsec.

Telescope back on M35, EL 63, AZ 258.

3:31 - We try to recenter the LGS within the FoV

New IM between LM1 and PatCams: 20140323_033220

We try to recenter also with a periscope loop

Shell RIP TN: 20140323_035507

JCG removes an actuator from the set. M2 restarted.

TN: 20140323_035833 Pupil loop closed

TN: 20140323_040024 Pupil & jitter loop closed

4:31 - We check the LAN flexures going from 45deg EL to Zentih: at each elevation we recenter the LGS with LM1-PatCams loop and we write down the LM1 motors positions:

EL A1 A2 A3
45 14470 8626 13265
50 14498 8649 13212
60 14577 8687 13095
70 14672 8710 12972
80 14802 8736 12818
90 14940 8739 12679
On the way back to EL45 we set on LM1 motors the previous values, then we close the LM1-PatCams loops and we write down the final position.

EL A1 A2 A3
80 14784 8741 12828
70 14672 8728 12962
60 14574 8695 13092
50 14498 8655 13207
45 14470 8637 13256
During the descending sweep the LGS are always found within the PatCams FoV using the ascending LM1 positions.

Another actuator dies and it is removed (560)

TN: 20140323_052208 EL45

TN: 20140323_052542 pupil loop closed, good for subapertures definition

TN: 20140323_052650 jitter loop closed

New slope_null acquired: 20140323_052924

We try to keep going the LM1-PatCams loop

We set a new PupilRecenteringTarget wrt the lab one: YELLOW = [+0.25; -0.25], RED = [+0.5; -0.5]. 20140323_054800

We redefine the subapertures wrt the lab ones: BLUE = [0; -1], YELLOW = [+3; -4], RED = [+5; -7]. New SubaperturesDefinition: 20140323_055200

New SlopeOffset: 20140323_055556

TN: 20140323_055940 half laser is closed

TN: 20140323_062326 good

LAN focus scan:

Stage [steps] TN Notes
0k 20140323_060444  
-10k 20140323_060518  
-20k 20140323_060546  
-30k 20140323_060643 FWHM: 2.2, 1.8
-40k 20140323_060902  
-60k 20140323_061010 FWHM: 1.95
-80k 20140323_061220 FWHM: 1.86
-100k 20140323_061309 FWHM: 1.98
-120k 20140323_061413 FWHM: 1.85
-140k 20140323_061535 FWHM: 1.92, 2.1, 1.86, 1.99
-160k 20140323_061657  
-180k 20140323_061852  
Sweep ToFfirst at big steps then at small ones:

ToF [us] TN
80 20140323_062611
82 20140323_062719
84 20140323_062758
86 20140323_062856
88 20140323_062939
84 20140323_063226
85 20140323_063318
86 20140323_063403
87 20140323_063455
The best position seems to be 86us

We set incremental modes on M2:

M2 Z5 [um] TN SlopeOffset
5 20140323_063851 20140323_055556
-5 20140323_064013 20140323_055556
-5 20140323_064222 20140323_064141
-2.5 20140323_064317 20140323_064141
-1.25 20140323_064408 20140323_064141
7:00 - Shell RIP, end of games

9:00 - Email to Armando

10:04 - Shell is back, Steve tries to move telescope to EL50, shell RIP again...

11:00 - Test of PI TF with sinusoidal command

Commanded freq [Hz] TN Notes
1 20140323_114547 A=1.35"
2 20140323_115229  
4 20140323_115753 wrong
4 20140323_120918 3.5Hz
8 20140323_121157 6.5Hz
16 20140323_121423 12Hz, A=1.3"
32 20140323_121620 20Hz, A=1.25"
64 20140323_121843 26Hz, A=1.25"
Test with a step command.

20140324

3:16 - M2 switched on, spotters are out

3:24 - dichroic is in

TN: 20140324_035339 we see clearly M2 spider. Good image to determine PatCams clock

LAS software dies, lasers off

4:55 - Too many clouds since ~1h

8:12 - Minaccia sereno! GR sends spotters out again

8:51 - Sweep ToF to check the clouds height. Nothing visible...
Topic revision: r9 - 04 Nov 2014, LorenzoBusoni
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback